OpenStudy (anonymous):

x^2 + 4x + 3 factor the solution?

OpenStudy (anonymous):

Want to know a trick?

OpenStudy (anonymous):

yes please.

OpenStudy (anonymous):

times the first X * 3 (last number) = 3.. What factors of 3 add up to 4?

OpenStudy (amistre64):

(m-2)^2 + 4(m-2) +3 = 0 where x = m-2 .... maybe; then it might be m+2, i aint determined that part yet :)

OpenStudy (anonymous):

3 and 1

OpenStudy (anonymous):

so your new middle numbers will be 3 and 1

OpenStudy (anonymous):

(x^2 +***3x***)(***1x***+3)

OpenStudy (anonymous):

So you factor out an X from the first one. x(x+3) and from the second one you can factor out a 1.. 1(x+3)

OpenStudy (amistre64):

m^2-4m+4 +4m-8 +3 ----------- m^2 -1 = 0 m = +-1 x = +-1 - 2 = -3,-1 dunno about the signs tho

OpenStudy (anonymous):

So you end up with x(x+3)+1(x+3) The x and the 1 on the outside is your first pair (x+1)

OpenStudy (anonymous):

then your second pair is (x+3)-- THEY ARE THE SAME

OpenStudy (anonymous):

So your answer would be (x+1)(x+3)

OpenStudy (amistre64):

woohoo; the roots are good :) (x+1)(x+3) lol

OpenStudy (anonymous):

Questions?

OpenStudy (anonymous):

You had to factor that one by PAIRING

OpenStudy (amistre64):

had to? nah

OpenStudy (anonymous):

Easier to understand.. Yes.

OpenStudy (anonymous):

:D

OpenStudy (amistre64):

perhaps :)